OpenStudy (j.f):

In the equations ax + by = 20 and bx + ay = 16, the value for x=2 and y=1. What are the coefficients a and b? 2a+b=20 2b+a=16 the coefficient of a and b is 2 and 1?

OpenStudy (anonymous):

put the numbers in

OpenStudy (anonymous):

the first tells you \[2a+b=20\] the second gives \[2b+a=16\]

OpenStudy (anonymous):

then either solve like any system \[2a+b=20\\ a+2b=16\] or solve say the first equation for \(b\) and substitute it in to the second one

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got this?

OpenStudy (j.f):

I understand you can solve the first and substitute that into the next...I just happen to forget how to solve the first

OpenStudy (misty1212):

\[2a+b=20\\ b=20-2a\]

OpenStudy (j.f):

yup I know that but 20 can't be subtracted by 2a because they don't share like variables